OpenStudy (anonymous):

My homework is about integers, and there's one part that says simplify.The questions are like 4x - 7x and -5x - 6x...I need help. :(

OpenStudy (anonymous):

4x-7x, they both have the same base and exponent (1), so you treat them as a normal number together. x(4-7)=-3x -5x-6x= ((-5)-6)x = -11x

OpenStudy (anonymous):

I also have questions like 10x - x and -x - (-15x) I don't know what to subtract since it's only an x.

OpenStudy (anonymous):

Pretend the x isn't there, and add it on afterwards. 10-1= 9(x) -1-(-15) = 14(x)